OpenStudy (teller):

Can someone help me please? The functions u and w are defined as follows. u(x)=2x+2 w(x)=-3x+4 Find the value of u(w(−5))

OpenStudy (anonymous):

Plug -5 into the equation for W, and then take whatever value you get from that and plug it into U. Does that make sense?

OpenStudy (teller):

So it would first look like this ? u(5)=2(5)+2

OpenStudy (teller):

I got u=12/5

OpenStudy (teller):

Not sure if that's correct or not.

OpenStudy (anonymous):

No. So what you have here is a composite function. Whenever you have a composite function, you always solve the function that's next to X first and work your way outward. So for example, if we have f(g(x)), we solve g(x) first, and then we solve f(x) with the number that we got from g(x).

OpenStudy (teller):

I'm confused..

OpenStudy (anonymous):

Plug -5 into W and tell me what you get

OpenStudy (teller):

- 11/5

OpenStudy (anonymous):

So -3(-5)+4 is -11/5? \[W(-5)=-3(-5)+4\]

OpenStudy (teller):

oh no -19/5

OpenStudy (anonymous):

The (-5) next do the W is not a specifically part of the equation. Whenever we have a function, let's say, F(x), we are saying that we have a function that is related in terms of x. Basically that further means that all of our variables in the equation are going to be in terms of X. For example, F(x)=x^2 or F(y)=y^2. It's not part of the equation, but it's a statement that helps clarify what variable our equation will use.

OpenStudy (teller):

I'm sorry I'm just confused now.....

OpenStudy (teller):

That confused me even more sorry....

OpenStudy (anonymous):

You're are dividing by -5 when you shouldn't be. It's not -5*W=-3*-5+4

OpenStudy (teller):

I'm not dividing I placed the equation like you wrote into the calculator and that's what I got...

OpenStudy (anonymous):

Yes you are. \[W(-5)=-3(-5)+4=19\] You said `-19/5` which is dividing by a -5.

OpenStudy (teller):

That's what I got when I placed it into the calculator... I placed the exact same problem you did what calculator are you using because I'm using mathway,,,

OpenStudy (anonymous):

That is because Mathway is interpreting it as "W times -5" and that's not what the question is asking you. It's asking you to evaluate the equation at -5.

OpenStudy (teller):

So what do I do then would the final answer be 53? Or -53

OpenStudy (anonymous):

\[W(-5)=-3(-5)+4=19\] \[U(19)=2(19)+2=40\]

OpenStudy (teller):

so it would be 40?

OpenStudy (anonymous):

Yes :)

OpenStudy (teller):

Thank you just math isn't my strongest subject just doesn't stick with me and my teacher knows this but yet she doesn't give a flying f about what happens .She just tells us to do it on our own and expect us to grasp it without lessons or even videos really.

OpenStudy (anonymous):

I can understand that frustration. Even professors do that in college and it really bothers me. I would recommend visiting her before or after class and see if she's willing to help. Sometimes teachers won't do much in class, but if you go to see them one-on-one and ask them for help, they'll be happy to. If she says she doesn't want to help you, then I would suggest going to administration or the principal to speak about it. And then find another math teacher before/after school to help you in the meantime.